Previous incidents
Tezos (Mainnet) - Outage
Resolved Aug 26 at 02:19pm UTC
The team found a routing issue and fixed it. The service has been reestablished.
We appreciate your patience and understanding.
1 previous update
Solana (Mainnet) - Service Degraded
Resolved Aug 23 at 12:08pm UTC
The service has been reestablished. Latency has returned to normal.
We appreciate your patience and understanding.
1 previous update
Near (Mainnet) - Service Degraded of historical archive data
Resolved Aug 21 at 11:51am UTC
The team has resolved the incident. We apologize for any inconvenience this may have caused and appreciate your patience during this time.
2 previous updates
REST API endpoints - Service Degraded
Resolved Aug 20 at 02:01pm UTC
The team has identified the problem and services are returning back to normal. We will be monitoring the situation further, but we are confident the issue has been resolved.
3 previous updates
Tron (Mainnet) - Service Degraded
Resolved Aug 15 at 03:33pm UTC
A HotFix has been deployed with additional mitigations.
We are observing an improvement in the overall performance.
We will keep closely monitoring the service during the next hours.
We appreciate your patience and understanding.
6 previous updates
Kaia (Mainnet) - Service Degraded
Resolved Aug 11 at 08:25am UTC
Regular operations have been restored. We appreciate your patience and understanding.
1 previous update
Cronos (Mainnet) - Outage
Resolved Aug 07 at 07:45am UTC
The team has restored access to the chain.
We appreciate your continued patience.
3 previous updates
Polygon (Mainnet) - Outage
Resolved Jul 30 at 12:06pm UTC
The service has been restored.
We appreciate your patience and understanding.
4 previous updates
REST API endpoints - Service Degraded
Resolved Jul 25 at 12:30pm UTC
We've observed sustained good performance since our previous update.
We appreciate your patience and understanding.
3 previous updates
Arbitrum One (Mainnet) - Service Degraded
Resolved Jul 25 at 07:56am UTC
We identified the issue with one of our providers and rerouted all traffic while they implemented a fix. The issue has now been resolved, and services are operating normally.
1 previous update
Kaia (Mainnet) - Outage
Resolved Jul 17 at 06:20am UTC
The team has reestablished the service.
We appreciate your patience and understanding.
1 previous update
Solana (Mainnet) - Partial Outage
Resolved Jul 14 at 07:40am UTC
The service has been reestablished.
We appreciate your patience and understanding.
1 previous update
ZCASH (Mainnet) - Service Degraded
Resolved Jul 16 at 06:51pm UTC
We have deployed zebrad nodes replacing Zcashd.
- https://zebra.zfnd.org/index.html
The chain performance has greatly improved.
2 previous updates
Algorand (Mainnet) - Service Degraded
Resolved Jul 11 at 11:32am UTC
The service has been reestablished.
We appreciate your patience and understanding.
2 previous updates
Solana (Mainnet) - Service Degraded
Resolved Jul 11 at 11:31am UTC
The service has been reestablished.
We appreciate your patience and understanding.
2 previous updates
Polygon (Mainnet) - Service Degraded
Resolved Jul 11 at 07:06pm UTC
This afternoon, the issues related to the Heimdall update on Polygon have stabilized.
Our team successfully restored the service, and we have been closely monitoring its performance, which has remained stable for the past few hours.
We appreciate your patience and understanding.
2 previous updates
Cosmos (Mainnet) - Outage
Resolved Jul 08 at 09:13am UTC
The team has recovered the service.
We appreciate your patience and understanding.
1 previous update
Solana (Mainnet) - Service Degraded
Resolved Jul 05 at 07:00pm UTC
The service has been reestablished.
We appreciate your patience and understanding.
3 previous updates
REST API endpoints - Service Degraded
Resolved Jul 02 at 02:46pm UTC
While we are awaiting a resolution from our hardware provider, our internal team has proactively rerouted and rebalanced services to restore availability.
At this time, the affected endpoints are operational and no further impact is expected.
We will continue to monitor the situation closely and follow up with the provider as needed.
2 previous updates
Polygon (Mainnet) - Service Degraded
Resolved Jul 01 at 07:04pm UTC
The latest Heimdall patch appears to have resolved the sync issues.
We appreciate your patience and understanding.
3 previous updates
Polygon Amoy (Testnet) - Service Degraded
Resolved Jul 04 at 12:47pm UTC
After the latest Erigon update, the nodes are following the tip again.
We appreciate your patience and understanding.
2 previous updates
Flare (Mainnet) - Service Degraded
Resolved Jun 26 at 10:58am UTC
The team has restored the service. Flare (Mainnet) nodes are following the tip again.
We appreciate your patience and understanding.
2 previous updates
Polygon Amoy (Testnet) - Outage
Resolved Jun 25 at 04:06pm UTC
The nodes were able to sync again with the tip after the Heimdall hard fork.
We appreciate your patience and understanding.
2 previous updates
Solana (Mainnet) - Service Outage
Resolved Jun 19 at 07:37am UTC
The team has recovered the service.
We appreciate your patience and understanding.
1 previous update
Polygon Amoy (Testnet) - Outage
Resolved Jun 14 at 09:03am UTC
Despite the issues with Erigon, The team has been able to sync the nodes with the tip.
We appreciate your patience and understanding.
1 previous update
Solana (Mainnet) - Service Degraded
Resolved Jun 13 at 03:57pm UTC
Further mitigations have been implemented. We have observed a stable latency for the past hour. We will keep monitoring the situation.
We appreciate your patience and understanding.
2 previous updates
XDC (Testnet) - Outage
Resolved Jun 12 at 05:30pm UTC
The service has been reestablished.
We appreciate your patience and understanding.
1 previous update
Polygon Amoy (Testnet) - Outage
Resolved Jun 12 at 07:53am UTC
We have reestablished service for Polygon Amoy (Testnet).
Our nodes are now following the chain tip.
We appreciate your patience and understanding.
1 previous update
Cosmos (Mainnet) - Upgrade v24.0.0
Resolved Jun 12 at 06:16am UTC
The nodes are in sync with the tip.
The service has been reestablished.
1 previous update
Solana (Mainnet) - Service Outage
Resolved Jun 09 at 07:45am UTC
The issue affecting Solana Mainnet nodes has been resolved.
All services are now operational. Our nodes are currently catching up to the latest block (tip) of the chain.
We are monitoring synchronization closely to ensure full stability and consistency across the network.
1 previous update
BSC (Testnet) - Service Degraded
Resolved Jun 06 at 08:44am UTC
The chain behaviour has returned to normal.
We will keep monitoring the situation.
1 previous update
Solana (Mainnet) - Service Degraded
Resolved Jun 03 at 04:31pm UTC
The performance issue has been resolved.
We appreciate your patience and understanding.
1 previous update